Circular No. 7560

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001G IN MCG +08-17-43 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong (Rolvenden, Kent) of an apparent supernova (mag 16.1) on a CCD frame taken on Jan. 8.051 UT with a 0.35-m Schmidt-Cassegrain telescope in the course of the U.K. Nova/Supernova Patrol. A confirming image obtained on Jan. 12.904 shows that the new object has brightened to mag ∼ 15.2. An average of measurements from each frame yields the following position for SN 2001G: α = 9^{h}09^{m}33^{s}.18, δ = +50^{o}16′51′′.3 (equinox 2000.0), which is 2′′ east and 4′′ south of the center of MCG +08-17-43. Nothing is present at this position to mag ∼ 18 on Armstrong’s patrol images from 2000 Nov. 30 and Dec. 27, neither is anything present on Palomar Sky Survey images obtained on 1991 Feb. 13 (red, limiting mag 20.8), 1996 Mar. 12 (infrared, limiting mag 19.5), and 1997 Feb. 5 (blue, limiting mag 22.5).

SUPERNOVA 2001H IN MCG –01-10-19 T. Puckett, Mountain Town, GA, provides his astrometry and pho-

tometry of an apparent supernova (mag ∼ 17.5) discovered by B. Holmes, Charleston, IL, on an unfiltered CCD frame taken with a 0.40-m reflector on Jan. 10.10 UT: α = 3^{h}40^{m}43^{s}.05, δ = −6^{o}24′59′′.1 (equinox 2000.0), which is 5′′.8 east and 2′′.9 south of the center of MCG –01-10-19. The new object was confirmed at mag 17.5 on an unfiltered CCD frame (limiting mag 20.5) taken by Puckett with the Puckett Observatory 0.60-m reflector on Jan. 13.01. SN 2001H does not appear on Palomar Sky Survey images (limiting mag ∼ 21.0) taken on 1987 Nov. 15 and 1982 Sept. 16.

COMETS C/1997 B3 AND C/1997 S3 (SOHO) Further to IAUC 7558, D. Hammer reports his measurements for ad-

ditional Kreutz sungrazing comets found by J. Danaher (Jan. comet in C3 data) and X. Leprette (Sept. comet in C2 data) on SOHO website images. The reduced observations and orbital elements by B. G. Marsden are given on the MPECs cited below.

Comet  1997 UT  α_{2000}  δ_{2000}  MPEC
C/1997 B3 Jan. 21.990  20^{h}35.^{m}8  −21^{o}27′  2001-A35
C/1997 S3 Sept.24.465  11 56.5  − 0 53  2001-A47
2001 January 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green